Solve the following equations. 
(8x - 3)/3x = 2 
Solve the following equations. 
(8x - 3)/3x = 2 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(8x-3)/3x=2
⇒8x/3x–3/3x=2
⇒8/3–1/x=2
⇒8/3–2=1/x
⇒(8-6)/3=1/x
⇒2/3=1/x
⇒x=3/2

Thanks